Demand for monopoly output at a price of 2, 3, 4, 5  is 80, 60, 50, 30.

 Determine:
1) max total revenue at the monopoly;
2) max total revenue if price discrimination will be used, for example, if  prices 3, 4, 5 will be set for different customer groups
1
Expert's answer
2021-01-08T12:58:49-0500

1) max total revenue at the monopoly is TR=4×50=200.

2) If price discrimination will be used, for example, if prices 3, 4, 5 will be set for different customer groups, then max total revenue is:

"TR = 3\u00d760 + 4\u00d750 + 5\u00d730 = 530."